# MegaScale-Infer: Serving Mixture-of-Experts at Scale with Disaggregated Expert Parallelism

### v3

@@ 6 Implementation
Fused kernels. To further improve efficiency and reduce latency, we implemented two types of fused kernels. The first is to overlap the communication of TP with the adjacent computation. Although intra-node TP typically uses high-speed interconnects like NVLINK for communication, it still introduces non-negligible overhead. To address this issue, we fuse its communication with the adjacent GEMM operation, such as implementing an all-gather and the following GEMM in a single kernel. The second is to fuse sequential memory-intensive operators. MoE includes several sequences of small memory-intensive operations. For example, attention nodes need to select top-k experts for each token after gating, compute intermediate results such as the number of tokens sent to each expert node and normalized token weights, and then perform data movement to scatter tokens to respective experts. We optimize this process by fusing these steps with the gating computation, reducing both kernel launch and memory access.
High-performance M2N communication library. We built our communication library as a Pytorch extension [15] in around 4900 and 5000 lines of C/C++ and Python code, respectively. As discussed earlier, our library is supported by technologies such as GPUDirect [13] and GDRCopy. We also carefully design network monitoring tools to delve into network and traffic-related optimizations.
@@ 7 Evaluation
In this section, we first evaluate the end-to-end performance of MegaScale-Infer against state-of-the-art LLM serving systems across various models and hardware configurations, including a heterogeneous environment. Next, we demonstrate the effectiveness of high-performance M2N communication through micro-benchmarks. Additionally, we conduct an ablation study to analyze the impact of ping-pong pipeline parallelism and deployment plans on MegaScale-Infer’s performance, highlighting the benefits of deployment plan optimization.
Model #Layers Hidden Size #Experts top-$k$ Intermediate Size
@@ 7.1 Experimental Setup
Testbed. We deploy MegaScale-Infer across two distinct clusters. The first cluster consists of eight nodes, each equipped with eight NVIDIA 80GB Ampere GPUs, 128 CPUs, 2 TB of host memory, and eight 200 Gbps Infiniband NICs. GPUs within the same node are interconnected via 400GB/s NVLINK. The second cluster is a heterogeneous setup, comprising two types of GPUs: NVIDIA H20 and NVIDIA L40S. H20 nodes are equipped with 900GB/s NVLINK and four 400 Gbps NICs, while L40S nodes utilize PCIe for intra-node communication and two 400 Gbps NICs for inter-node communication. As shown in Table 3, H20 also offers higher memory capacity and bandwidth but has less computational power than L40S.
@@
Models and workload. We evaluate MegaScale-Infer using Mixtral 8x22B [63], DBRX [64], and a scaled MoE model, which shares a similar structure but includes more experts. They contain 141B, 132B, and 317B parameters, respectively. The model configurations are detailed in Table 4. For all experiments, the data type for weights, activations, and KV cache are bfloat16. We obtain a real-world dataset from our production and use it as the experimental workload. The median input and output length are 571 and 159 tokens, respectively.
@@
Baselines. We compare MegaScale-Infer with two state-of-the-art serving systems: vLLM [44] (version 0.6.6.post1) and TensorRT-LLM [51] (version 0.15.0). Both systems support popular techniques for LLM serving, including FlashAttention [27], PagedAttention [44], and continuous batching [72]. They primarily rely on tensor parallelism for distributed LLM serving, with TensorRT-LLM additionally supporting expert parallelism for expert layers. For larger models requiring inter-node communication, both systems also support pipeline parallelism. Due to GPU memory limits and large model sizes, serving Mixtral 8x22B and DBRX with these baselines requires a minimum of 8 GPUs, while the scaled-MoE necessitates multi-node deployment.
Metrics. We focus on maximizing decoding throughput normalized by cost while adhering to a time per output token (TPOT) latency requirement. Specifically, for homogeneous deployment, we use per-GPU decoding throughput, i.e., tokens generated per second divided by the number of GPUs, as the primary metric. For heterogeneous deployment, we report per-cost decoding throughput. Following prior work [74], we set the TPOT requirement to 150 milliseconds. Additionally, we present the median and tail latency and throughput of M2N communication.
@@ 7.2 End-to-end Experiment
Figure 8: Normalized decoding throughput of different models on NVIDIA Ampere GPUs.
@@
Homogeneous deployment. We first evaluate the performance of MegaScale-Infer with different MoE models on NVIDIA 80GB Ampere GPUs. Both vLLM and TensorRT-LLM serve Mixtral 8x22B and DBRX on a single node and serve Scaled-MoE across two nodes. The results are presented in Figure 8. Since vLLM deploys and serves the model as a whole, the batch size for the expert modules tends to be small, resulting in low GPU utilization. TensorRT-LLM achieves higher throughput than vLLM through custom kernel optimizations, but it also adopts a holistic service approach, which means it cannot avoid the issue of low GPU utilization in the FFN modules. By separating the attention and FFN modules, MegaScale-Infer aggregates batched requests from multiple attention modules, which increases the FFN batch size. This shift helps transition FFN from being memory-intensive to compute-intensive, thereby improving GPU utilization. As a result, MegaScale-Infer achieves 2.56×2.56$\times$2.56 × and 1.28×1.28$\times$1.28 × higher per-GPU decoding throughput than vLLM and TensorRT-LLM. For Scaled-MoE, the expensive inter-node communication overhead, coupled with certain implementation limitations in a multi-node environment, results in even lower GPU utilization for the baselines. In contrast, MegaScale-Infer is deployed across multiple nodes for all models due to its disaggregated deployment, but it overlaps computation and communication through the design of ping-pong pipeline parallelism. Consequently, MegaScale-Infer improves the decoding throughput per GPU for Scaled-MoE by 7.11×7.$1$ and 1.90×1.90$\times$1.90 × compared to vLLM and TensorRT-LLM, respectively.
Figure 9: Normalized decoding throughput (throughput divided by normalized cost) of different models on NVIDIA H20 and L40S GPUs.
@@
Heterogeneous deployment. To demonstrate the benefits of MegaScale-Infer under heterogeneous deployment, we build a cluster consisting of NVIDIA H20 and L40S GPUs and conduct experiments on it. Since neither vLLM nor TensorRT-LLM supports heterogeneous deployment, we separately evaluate them on H20 and L40S. To fully leverage the capacity of each GPU type, MegaScale-Infer assigns H20 for attention modules and L40S for experts. Figure 9 presents the performance measured by decoding throughput per cost. Here we define cost with the normalized purchase price as shown in Table 3, which can easily be replaced by the rental price for cloud service users. Due to the 48GB memory capacity of L40S, all models in both baselines require a multi-node setup. In addition, the relatively weak intra-node and inter-node communication performance of the L40S leads to low GPU utilization for both vLLM and TensorRT-LLM. In contrast, H20 is more suitable for LLM serving due to its large memory capacity, higher bandwidth, and faster communication. As a result, vLLM and TensorRT-LLM achieve higher decoding throughput on H20. However, the L40S also offers unique advantages, particularly its high compute power per unit cost, making it well-suited for executing compute-intensive tasks. Our heterogeneous deployment simultaneously maximizes the advantages of the high bandwidth of H20 and the cost-effective compute power of L40S. This results in an improvement of up to 3.24×3.24$\times$3.24 × and 1.86×1.86$\times$1.86 × on the unit cost throughput compared to vLLM and TensorRT-LLM on H20, respectively.
@@ 7.4 Ablation Study
Figure 13: Latency and throughput of DBRX under different DP degree of attention. (a) Per output token latency. (b) Decoding throughput normalized by number of GPUs.
@@ 8 Related Work
Resource disaggregation. Disaggregating hardware resources into separate resource pools allows for independent scaling, resulting in more efficient deployments. Several systems [33, 57] adopt this approach. In the context of LLM serving, the distinct characteristics of the prefill and decoding phases make their disaggregation a widely used solution [74, 53, 37, 60, 54]. MegaScale-Infer also employs this approach and further optimizes MoE serving efficiency during decoding by disaggregating the attention and FFN modules. FASTDECODE [35], Lamina [25], and MoE-Lightning [23] offload attention computation to the cheaper devices, such as CPU, during decoding. However, offloading to those devices results in slow performance, and the challenges posed by MoE’s sparsity remain unresolved.
@@
Collective communication for ML. Distributed machine learning jobs heavily rely on high-performance collective communications, such as all-reduce and all-to-all, to achieve high throughput and low latency. NVIDIA NCCL [6] is the most popular collective communication library in both industry and academia. SCCL [22], TACCL [56] and TE-CCL [50] propose automatic synthesis of optimal collective communication algorithms tailored to distinct hardware and topologies. CoCoNET [41] and Centauri [24] improve performance by overlapping communication with computation in distributed machine learning. The disaggregation of attention and FFN in MoE necessitates a new form of collective communication, i.e., M2N. We identify and eliminate the overhead and instability present in existing solutions.

### v4

@@ 6 Implementation
Fused kernels. To further improve efficiency and reduce latency, we implemented two types of fused kernels. The first one is to overlap the communication of TP with the adjacent computation. Although intra-node TP typically uses high-speed interconnects like NVLINK for communication, it still introduces non-negligible overhead. To address this issue, we utilize Flux [29] to fuse communication with the adjacent GEMM operation, such as implementing an all-gather and the following GEMM in a single kernel. The second one is to fuse sequential memory-intensive operators. MoE includes several sequences of small memory-intensive operations. For example, attention nodes need to select top-k experts for each token after gating, compute intermediate results such as the number of tokens sent to each expert node and normalized token weights, and then perform data movement to scatter tokens to respective experts. We optimize this process by fusing these steps with the gating computation, reducing both kernel launch and memory access.
High-performance M2N communication library. We built our communication library as a Pytorch extension [21] in around 4900 and 5000 lines of C/C++ and Python code, respectively. Our library is supported by technologies such as GPUDirect [17] and GDRCopy. We also carefully design network monitoring tools to delve into network and traffic-related optimizations.
@@ 7 Evaluation
[…]
@@
Figure 8: Performance of MegaScale-Infer on NVIDIA Ampere GPUs.
In this section, we first evaluate the end-to-end performance of MegaScale-Infer against state-of-the-art LLM serving systems across various models and hardware configurations, including a heterogeneous environment. Next, we demonstrate the effectiveness of high-performance M2N communication through micro-benchmarks. Additionally, we conduct an ablation study to analyze the impact of disaggregated architecture and M2N optimization on MegaScale-Infer’s performance. We also present the effectiveness of ping-pong pipeline parallelism and deployment plans, highlighting the benefits of deployment plan optimization.
@@ 7.1 Experimental Setup
Testbed. We deploy MegaScale-Infer across two distinct clusters. The first cluster consists of eight nodes, each equipped with eight NVIDIA 80GB Ampere GPUs, 128 CPUs, 2 TB of host memory, and eight 200 Gbps Infiniband NICs. GPUs within the same node are interconnected via 400GB/s NVLINK. The second cluster is a heterogeneous setup, comprising two types of GPUs: NVIDIA H20 and L40S. H20 nodes are equipped with 900GB/s NVLINK and four 400 Gbps NICs, while L40S nodes utilize PCIe for intra-node communication and two 400 Gbps NICs for inter-node communication. As shown in Table 3, H20 offers higher memory capacity and bandwidth but has less computational power than L40S.
@@
Models and workload. We evaluate MegaScale-Infer using Mixtral 8x22B [70], DBRX [71], and a Scaled-MoE, which shares a similar structure but includes more experts. They contain 141B, 132B, and 317B parameters, respectively. The model configurations are detailed in Table 4. For all experiments, the data types for weights, activations, and KV cache are bfloat16. We obtain a dataset from our production and use it as the experimental workload. The median input and output length are 571 and 159 tokens, respectively.
@@
Baselines. We compare MegaScale-Infer with two state-of-the-art serving systems: vLLM [51] and TensorRT-LLM [57]. Both systems support popular techniques for LLM serving, including FlashAttention [33], PagedAttention [51], and continuous batching [79]. They primarily rely on tensor parallelism for distributed LLM serving, with TensorRT-LLM additionally supporting expert parallelism for expert layers. For larger models requiring inter-node communication, both systems also support pipeline parallelism. Due to GPU memory limits and large model sizes, serving Mixtral 8x22B and DBRX with these baselines requires a minimum of 8 GPUs, while Scaled-MoE necessitates multi-node deployment. The Attention-FFN disaggregation architecture within MegaScale-Infer naturally adapts to prefill/decoding (P/D) disaggregation, effectively preventing interference between these two phases. Existing baselines are still in the process of supporting or optimizing P/D disaggregation. To ensure a fair comparison and to accurately quantify the performance gains under P/D disaggregation, we evaluate all baselines and MegaScale-Infer by temporally separating their prefill and decoding phases. Specifically, when measuring decoding throughput and time between tokens, we exclude the prefill phase and consider only the iteration time and output tokens of the decoding phase.
Metrics. The primary objective of our work is to improve the efficiency of MoE inference during the decoding phase, which suffers from low GPU utilization due to its memory-bandwidth-bound nature. Enhancing decoding efficiency in MoE inference is the key to reducing overall serving costs. Therefore, we focus on maximizing the cost-normalized decoding throughput, subject to a specified time-between-tokens (TBT) latency constraint. Specifically, for homogeneous deployment, we use per-GPU decoding throughput, i.e., tokens generated per second, excluding the first output token, divided by the number of GPUs, as the primary metric. For heterogeneous deployment, we report per-cost decoding throughput. Following prior work [82], we set the TBT requirement to 150 milliseconds. We also report the mean time between tokens corresponding to the throughput results. While our primary contribution lies in improving the efficiency of the decoding phase, we also report end-to-end throughput results that include the generation of the first output token. This offers a more comprehensive evaluation of MegaScale-Infer’s performance improvements across both the prefill and decoding phases. Notably, we observe that heterogeneous deployment also yields benefits for the prefill phase. We present the throughput per unit power under heterogeneous deployment. Additionally, we present the median and tail latency and throughput of M2N communication.
Figure 9: Performance of MegaScale-Infer on NVIDIA H20 and L40S GPUs.
@@ 7.2 End-to-end Experiment
Homogeneous deployment. We first evaluate the performance of MegaScale-Infer with different MoE models on NVIDIA 80GB Ampere GPUs. Both vLLM and TensorRT-LLM serve Mixtral 8x22B and DBRX on a single node and serve Scaled-MoE across two nodes. The results are presented in Figure 8. Since vLLM deploys and serves the model as a whole, the batch size for the expert modules tends to be small, resulting in low GPU utilization. TensorRT-LLM achieves higher throughput than vLLM through custom kernel optimizations, but it also adopts a holistic service approach, which means it cannot avoid the issue of low GPU utilization in the FFN modules. By separating the attention and FFN modules, MegaScale-Infer aggregates batched requests from multiple attention modules, which increases the FFN batch size. This shift helps transition FFN from being memory-intensive to compute-intensive, thereby improving GPU utilization. As a result, MegaScale-Infer achieves 2.56×2.56$\times$2.56 × and 1.28×1.28$\times$1.28 × higher per-GPU decoding throughput than vLLM and TensorRT-LLM, as shown in Figure 8(a). For Scaled-MoE, the expensive inter-node communication overhead, coupled with certain implementation limitations in a multi-node environment, results in even lower GPU utilization for the baselines. In contrast, MegaScale-Infer is deployed across multiple nodes for all models due to its disaggregated deployment, but it overlaps computation and communication through the design of ping-pong pipeline parallelism. Consequently, MegaScale-Infer improves the decoding throughput per GPU for Scaled-MoE by 7.11×7.$1$ and 1.90×1.90$\times$1.90 × compared to vLLM and TensorRT-LLM, respectively.
Figure 8(b) shows the mean time between tokens of MegaScale-Infer and the baselines corresponding to the decoding throughput in Figure 8(a). Due to the disaggregation of attention and expert modules, MegaScale-Infer introduces cross-node communication at every layer, which affects latency. Ping-pong pipeline parallelism, while overlapping communication with computation across micro-batches, does not reduce the per-token latency for an individual micro-batch. Aiming for full GPU utilization, this approach may even incur additional latency, as indicated by constraint 3. Nevertheless, MegaScale-Infer effectively mitigates the communication overhead by employing a high-performance M2N communication library. As a result, MegaScale-Infer achieves a time between tokens comparable to those of the baseline systems.
Figure 8(c) shows the end-to-end per-GPU throughput including the prefill phase. As the prefill phase is predominantly compute-bound, our approach does not yield performance improvements for this stage under homogeneous deployments. Consequently, when the prefill phase is taken into account, the overall end-to-end performance gain is less pronounced compared to the decoding phase alone. Nevertheless, MegaScale-Infer still achieves up to a 1.18$\times$ improvement in throughput, demonstrating its effectiveness in end-to-end scenarios.
Heterogeneous deployment. To demonstrate the benefits of MegaScale-Infer under heterogeneous deployment, we build a cluster consisting of NVIDIA H20 and L40S GPUs and conduct experiments on it. Since neither vLLM nor TensorRT-LLM supports heterogeneous deployment, we separately evaluate them on H20 and L40S. To fully leverage the capacity of each GPU type, MegaScale-Infer assigns H20 for attention modules and L40S for experts. Figure 9(a) presents the performance measured by decoding throughput per unit cost. Here we define cost with the normalized purchase price as shown in Table 3, which can easily be replaced by the rental price for cloud service users. Due to the 48GB memory capacity of L40S, all models in both baselines require a multi-node setup. In addition, the relatively weak intra-node and inter-node communication performance of the L40S leads to low GPU utilization for both vLLM and TensorRT-LLM. In contrast, H20 is more suitable for LLM serving due to its large memory capacity, higher bandwidth, and faster communication. As a result, vLLM and TensorRT-LLM achieve higher decoding throughput on H20. However, the L40S also offers unique advantages, particularly its high compute power per unit cost, making it well-suited for executing compute-intensive tasks. Our heterogeneous deployment simultaneously maximizes the advantages of the high bandwidth of H20 and the cost-effective compute power of L40S. This results in an improvement of up to 3.24×3.24$\times$3.24 × and 1.86×1.86$\times$1.86 × on the unit cost decoding throughput compared to vLLM and TensorRT-LLM on H20, respectively.
Figure 10: Throughput per unit power of MegaScale-Infer on NVIDIA H20 and L40S GPUs.
@@
The latency results corresponding to Figure 9(a) are shown in Figure 9(b). Similar to the homogeneous deployment scenario, the mean time between tokens under heterogeneous deployment remains comparable to those of the baselines. Moreover, when compared to the baselines deployed exclusively on L40S GPUs, our approach achieves slightly improved latency performance.
We further apply heterogeneous deployment to the prefill phase and report the end-to-end throughput results, including prefill computation, in Figure 9(c). While heterogeneous deployment does not enhance resource utilization during the prefill phase, it effectively reduces inference costs by offloading expert computations to the more cost-efficient L40S GPUs. As a result, when evaluating end-to-end performance across both the prefill and decoding phases, MegaScale-Infer achieves up to a 1.66$\times$ improvement in throughput per unit cost compared to the baselines. In summary, MegaScale-Infer is particularly well-suited for heterogeneous deployment, as it significantly lowers inference costs across the entire serving pipeline.
We also evaluate the impact of heterogeneous deployment on power, and the results are presented in Figure 10. Since the H20 and L40S GPUs each offer lower energy consumption per unit of bandwidth and compute, respectively, heterogeneous deployment across these two GPU types can also improve throughput per unit power. This improvement is observed in both the decoding and prefill phases. As shown in Figure 10(a) and (b), our system achieves 1.80$\times$ and 1.72$\times$ higher decoding and end-to-end throughput per unit power, respectively, compared to the baseline.
@@ 7.4 Ablation Study
Effectiveness of disaggregated expert parallelism and M2N optimization. Figure 13 illustrates the performance gains achieved on Ampere GPUs through disaggregated expert parallelism and M2N communication optimizations. We select vLLM as the baseline, as it colocates attention and expert modules, and its kernel performance—particularly for matrix multiplication and attention operations—is comparable to our implementation. By adopting a disaggregated architecture, requests from multiple attention modules can be aggregated, thereby increasing the effective batch size on the expert side and improving the computational efficiency of expert modules. As a result, even when using NCCL as the backend for M2N communication, the disaggregated approach achieves up to a 4.66$\times$ throughput improvement over the colocated baseline. By leveraging our optimized M2N communication library, we further reduce communication overhead, enabling the M2N communication time of a single micro-batch to fall below its computation time (satisfying constraint 2). As a result, communication can be fully overlapped with computation through the use of the ping-pong pipeline, leading to an additional throughput improvement of up to 1.53$\times$.
Figure 13: Effectiveness of disaggregated expert parallelism and M2N optimization.
@@
Figure 15: Performance of DBRX under different DP degree of attention. (a) Per output token latency. (b) Decoding throughput normalized by number of GPUs.
8 Deployment Experience
[] Refer to caption [] Refer to caption [] Refer to caption
Figure 16: Expert load distribution in real traffic. (a) Received token count of each expert in a batch. (b) Received token ratio of each expert during decoding. (c) Received token ratio of each expert during prefill.
MegaScale-Infer has been deployed in the company’s production inference services and is operating on a cluster with nearly 10,000 GPUs. Under heterogeneous deployment, it reduces the cost of serving the same traffic by 1.5–2.0$\times$, depending on the workload characteristics.
Expert balance. In real-world deployment, we gained additional insights into expert load distribution. Figure 16 illustrates the number of tokens processed by each expert across four model layers for a single batch. There is a significant imbalance in the load distribution, with some experts being substantially hotter than others. Further analysis of different phases reveals additional patterns. As shown in Figures 16 and 16, we sample four batches executed within a short time window and measure the load on all experts in a specific layer. During the decoding phase, expert load remains relatively stable across batches, whereas in the prefill phase, it fluctuates more significantly. These observations motivate us to adopt a static or periodic expert load balancing strategy during decoding, while employing more frequent expert plan adjustments in the prefill phase.
Attention balance. We also observe load imbalance on the attention side. Due to variations in sequence lengths, attention nodes can experience significantly different computation times even when processing batches of the same size. Since MegaScale-Infer performs frequent synchronization across all nodes, such imbalances can introduce bubbles and degrade overall system efficiency. To mitigate this issue, we profile the runtime of key operators under varying sequence lengths and batch sizes to estimate the computation time of a given batch. We then compose the batch on each attention node to match a predefined target execution time, thereby balancing the workload across nodes.
@@ 9 Related Work
Resource disaggregation. Disaggregating hardware resources into separate resource pools allows for independent scaling, resulting in more efficient deployments. Several systems [40, 64] adopt this approach. In the context of LLM serving, the distinct characteristics of the prefill and decoding phases make their disaggregation a widely used solution [82, 60, 44, 67, 61]. MegaScale-Infer also employs this approach and further optimizes MoE serving efficiency during decoding by disaggregating attention and FFN modules. FASTDECODE [42], Lamina [31], and MoE-Lightning [28] offload attention computation to cheaper devices, such as CPU, during decoding. However, offloading results in higher latency, and the challenges posed by MoE’s sparsity remain unresolved.
@@
Collective communication for ML. Distributed machine learning jobs heavily rely on high-performance collective communications, such as all-reduce and all-to-all, to achieve high throughput and low latency. NVIDIA NCCL [8] is the most popular collective communication library in both industry and academia. SCCL [27], TACCL [63], and TE-CCL [56] propose automatic synthesis of optimal collective communication algorithms tailored to distinct hardware and topologies. CoCoNET [48] and Centauri [30] improve performance by overlapping communication with computation in distributed machine learning. The disaggregation of attention and FFN in MoE necessitates a new form of collective communication, i.e., M2N. We identify and eliminate the overhead and instability present in existing solutions.